# GASE: Generatively Augmented Sentence Encoding

### v1

@@ Abstract
We propose an approach to enhance sentence embeddings by applying generative text models for data augmentation at inference time. Unlike conventional data augmentation that utilises synthetic training data, our approach does not require access to model parameters or the computational resources typically required for fine-tuning state-of-the-art models. Generatively Augmented Sentence Encoding uses diverse linguistic synthetic variants of input texts generated by paraphrasing, summarising, or extracting keywords, followed by pooling the original and synthetic embeddings. Experimental results on the Massive Text Embedding Benchmark for Semantic Textual Similarity (STS) demonstrate performance improvements across a range of embedding models using different generative models for augmentation. We find that generative augmentation leads to larger performance improvements for embedding models with lower baseline performance. These findings suggest that integrating generative augmentation at inference time adds semantic diversity and can enhance the robustness and generalizability of sentence embeddings for embedding models. Our results show that the degree to which generative augmentation can improve STS performance depends not only on the embedding model but also on the dataset. From a broader perspective, the approach allows trading training for inference compute.
@@
Manuel Frank mfr-research@protonmail.com                        Haithem Afli Department of Computer Science Munster Technological University Haithem.Afli@mtu.ie
@@ 1 Introduction
Representation learning has emerged as a fundamental technique in natural language processing (NLP). Still, the quality and robustness of embeddings largely depend on the richness and diversity of the training data. To derive sentence embeddings from BERT (Devlin et al., 2019) word embeddings, these can be averaged or the embedding of the CLS token [^1] CLS tokens are special tokens placed at the beginning of each input example in a BERT model, providing a representation of the entire input for use in classification tasks (Devlin et al., 2019). being used. Based on encoder-style transformers, such as BERT, Sentence Transformers (STs) (Reimers and Gurevych, 2020) were trained to directly output sentence embeddings by calculating an attention-weighted average of the word embeddings. These sentence embeddings can be used in pairwise tasks, e.g. paraphrase identification, or non-pairwise tasks, e.g. sentiment analysis. Recent advancements in generative models, such as OpenAI’s GPT-4 (OpenAI, 2023), Anthropic’s Claude 3 models (Anthropic, 2024) and Google’s Gemini models (Gemini Team et al., 2024), have shown remarkable capabilities in generating human-like text. Moreover, generative models like ChatGPT have been demonstrated to be capable of paraphrasing, e.g. for data augmentation Dai et al. (2023). The authors use ChatGPT to rephrase a given text to generate synthetic data which is used to train a downstream BERT model generating embeddings for text classification. To augment generative models, the integration with representation learning has been explored, e.g. in Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020). In contrast, we introduce Generatively Augmented Sentence Encoding (GASE), an approach to augment embedding models using generative Large Language Models (LLMs). Instead of generating synthetic data for pre-training or fine-tuning, we propose an augmentation approach which is applied at inference time by generating one or more textual variants of input data and pooling the embeddings of the original input text and the synthetic texts. The underlying hypothesis of our work on generative augmentation is that generative models can introduce semantic diversity which is not modelled by current embedding models without generative augmentation and, therefore, can benefit the performance of downstream tasks using the embeddings. Our method involves three generative tasks to generate the synthetic data from the original data: paraphrasing, summarising, and extracting keywords. To the best of our knowledge, this is the first work proposing generative models to augment sentence encoders at inference time.
2 Approach
As illustrated in Figure 1, GASE consists of the following steps:
1. Generative Augmentation. In this step $k$ generative text models are used to produce $k$ variations of the original input sequence. We propose three augmentation functions: Paraphrasing, which provides a semantically equivalent but lexically or syntactically different variation of the input text. Summarising which produces a shorter output text that captures the most important information of the input text. And Extracting Keywords which lists the most relevant words from a given text. An example for each is provided in Figure 2.
@@ 2 Approach
We evaluate $k\in\{0,1,2\}$ using the generative LLMs GPT-3.5 Turbo (OpenAI, 2024b), Reka-Flash (Reka, 2024), and GPT-4o mini  (OpenAI, 2024a) with temperature and top_p set to 00 and $1$ respectively. As a baseline, we use $k=0$, i.e., no augmentation. The prompts can be found in Appendix A. Model-specific post-processing is performed to remove meta-text that is not used for augmentation.
2. Sentence Embeddings. For the original sequence and each of the $k$ synthetic variations an embedding is inferred. We evaluated the following embedding models which include non-transformer baselines, locally run ST models, and SOTA models accessed via an API using default hyperparameter values: glove.840B.300d (Pennington et al., 2014), bert-large-cased (Devlin et al., 2019), all-MiniLM-L6-v2 Sentence Transformers (2024), all-mpnet-base-v2 Sentence Transformers (2024), embed-english-light-v3.0 Cohere (2024), embed-english-v3.0 Cohere (2024), voyage-2 (Voyage AI, 2024), voyage-lite-02-instruct,  (Voyage AI, 2024), voyage-large-2 (Voyage AI, 2024), voyage-large-2-instruct (Voyage AI, 2024), mxbai-embed-large-v1 Mixedbread (2024).
3. Pooling. The embeddings of the original text and the $k$ augmented texts are pooled by computing their arithmetic mean.
We evaluate our approach on the English subtasks of the MTEB STS task [^2] This includes STSBenchmark (Cer et al., 2017a), STS12 (Agirre et al., 2012), STS13 (Agirre et al., 2013), STS14 Agirre et al. (2014), STS15 (Agirre et al., 2015), STS16 (Agirre et al., 2016), STS17 (Cer et al., 2017b), STS22 (Chen et al., 2022), SICK-R Marelli et al. (2014), and BIOSSES (Soğancıoğlu et al., 2017). According to the original paper (Muennighoff et al., 2023), MTEB also includes the STS11 dataset. However, the MTEB leaderboard (Hugging Face, 2024) does not. To make our results comparable to the leaderboard we excluded STS11. using cosine similarity and Spearman’s rank correlation [^3] See  Reimers et al. (2016) for why Spearman’s rank correlation is preferable over Pearson correlation for STS tasks..
3 Results
Table 1 shows the results for different augmentation strategies using GPT-3.5 Turbo vs. the respective baseline. (See Table 4 in the Appendix for the results per dataset.) In terms of average scores across datasets, all embedding models benefit from some form of augmentation.
Embedding
model
Augmentation
strategy
Avg.    STS22
glove. 840B.300d - 57.86    54.08
paraphrasing 59.98    57.78
summarising 60.64    61.21
extracting keywords 60.82    56.94
bert- large-cased - 60.78    55.48
paraphrasing 66.29    58.57
summarising 66.89    63.47
extracting keywords 62.62    54.64
all-MiniLM- L6-v2 - 78.91    67.26
paraphrasing 80.32    68.47
summarising 79.90    69.27
extracting keywords 77.96    64.57
all-mpnet- base-v2 - 80.28    68.00
paraphrasing 81.39    69.01
summarising 81.48    70.12
extracting keywords 79.95    66.19
embed-english- light-v3.0 - 78.75    67.88
paraphrasing 80.30    68.42
summarising 80.21    68.97
extracting keywords 78.53    67.27
embed- english-v3.0 - 81.25    68.22
paraphrasing 82.10    68.63
summarising 82.08    70.29
extracting keywords 81.26    66.91
voyage-2 - 82.50    65.26
paraphrasing 83.38    66.80
summarising 82.74    68.50
extracting keywords 81.21    65.29
voyage-lite- 02-instruct - 85.95    78.62
paraphrasing 85.99    79.08
summarising 85.43    79.87
extracting keywords 83.91    76.60
voyage- large-2 - 83.60    63.97
paraphrasing 83.92    64.69
summarising 83.66    65.90
extracting keywords 81.98    63.68
voyage-large- 2-instruct - 84.61    66.59
paraphrasing 84.80    67.72
summarising 84.54    68.66
extracting keywords 83.40    65.73
mxbai-embed- large-v1 - 84.63    68.73
paraphrasing 85.01    70.03
summarising 84.45    70.36
extracting keywords 83.42    67.83
Table 1: Average scores on the MTEB datasets and STS22 individually for all embedding models with and without generative augmentation using GPT-3.5 Turbo (Spearman’s rank correlation in %, bold: highest scores).
The largest performance increases are observed for BERT (+$6.11$pp) and GloVe (+$2.96$pp). GloVe is also the only embedding model achieving its best results using keyword extraction. Moreover, BERT and MPNET STs benefit most from summarisation. In terms of datasets, all embedding models scored lowest on STS22.
@@ 3 Results
Table 2: Mean word count (using GPT-3.5 Turbo for generative augmentation).
As expected, summarising and extracting keywords significantly reduced the mean word count on average across datasets (see Table 2). While paraphrases for all datasets except STS22 are close to the word count of the original texts, GPT-3.5 Turbo does not maintain STS22’s high word count.
@@
Since paraphrase augmentation provided the largest improvements for most models, we evaluated it for all generative models (see Table 3 and for results per dataset Table 5 in the Appendix). Using a single generative model, GPT-3.5 Turbo achieved the best results for all embedding models except GloVe and Voyage-Large-2 which performed better using Reka-Flash. Hence, we ran experiments with these two models generating $k=2$ paraphrases.
Across embedding models, the use of these two generative models led to higher scores than a single generative model. The largest performance gains were achieved with BERT (+$7.23$pp) and GloVe (+$4.11$pp). MiniLM ST, MPNet ST, Cohere’s Embed models and Voyage-2 demonstrated maximum improvements between $1.28$ and $1.79$pp while the remaining embedding models showed performance gains below $1$pp. Notably, the results with $k=2$ surpass the mean between the scores of employing GPT-3.5 Turbo and Reka Flash individually.
@@
Generative model
Avgerage
score
glove. 840B.300d - 57.86
gpt-3.5-turbo 59.98
reka-flash 60.73
GPT-4o mini 60.04
gpt-3.5-turbo, reka-flash 61.97
bert- large-cased - 60.78
gpt-3.5-turbo 66.29
reka-flash 65.76
GPT-4o mini 65.01
gpt-3.5-turbo, reka-flash 68.01
all-MiniLM- L6-v2 - 78.91
gpt-3.5-turbo 80.32
reka-flash 79.74
GPT-4o mini 79.81
gpt-3.5-turbo, reka-flash 80.64
all-mpnet- base-v2 - 80.28
gpt-3.5-turbo 81.39
reka-flash 81.01
GPT-4o mini 81.07
gpt-3.5-turbo, reka-flash 81.72
embed-english- light-v3.0 - 78.75
gpt-3.5-turbo 80.30
reka-flash 79.72
GPT-4o mini 79.90
gpt-3.5-turbo, reka-flash 80.54
embed- english-v3.0 - 81.25
gpt-3.5-turbo 82.10
reka-flash 82.00
GPT-4o mini 81.82
gpt-3.5-turbo, reka-flash 82.53
voyage-2 - 82.50
gpt-3.5-turbo 83.38
reka-flash 83.37
GPT-4o mini 83.07
gpt-3.5-turbo, reka-flash 83.80
voyage-lite- 02-instruct - 85.95
gpt-3.5-turbo 85.99
reka-flash 85.92
GPT-4o mini 85.97
gpt-3.5-turbo, reka-flash 86.05
voyage- large-2 - 83.60
gpt-3.5-turbo 83.92
reka-flash 83.98
GPT-4o mini 83.95
gpt-3.5-turbo, reka-flash 84.12
voyage-large- 2-instruct - 84.61
gpt-3.5-turbo 84.80
reka-flash 84.68
GPT-4o mini 84.78
gpt-3.5-turbo, reka-flash 84.86
mxbai-embed- large-v1 - 84.63
gpt-3.5-turbo 85.01
reka-flash 84.80
GPT-4o mini 84.90
gpt-3.5-turbo, reka-flash 85.17
Table 3: Average scores on the MTEB datasets with and without paraphrase augmentation using GPT-3.5 Turbo, GPT-4o mini, and Reka-Flash (Spearman’s rank correlation in %, bold: highest scores).
Figure 3: Score without augmentation vs. max. score with augmentation per embedding model.

### v2

@@ Abstract
We propose a training-free approach to improve sentence embeddings leveraging test-time compute by applying generative text models for data augmentation at inference time. Unlike conventional data augmentation that utilises synthetic training data, our approach does not require access to model parameters or the computational resources typically required for fine-tuning state-of-the-art models. Generatively Augmented Sentence Encoding variates the input text by paraphrasing, summarising, or extracting keywords, followed by pooling the original and synthetic embeddings. Experimental results on the Massive Text Embedding Benchmark for Semantic Textual Similarity (STS) demonstrate performance improvements across a range of embedding models using different generative models for augmentation. We find that generative augmentation leads to larger performance improvements for embedding models with lower baseline performance. These findings suggest that integrating generative augmentation at inference time adds semantic diversity and can enhance the robustness and generalisability of sentence embeddings for embedding models. Our results show that performance gains depend on the embedding model and the dataset.
@@
Manuel Frank Department of Computer Science Munster Technological University Manuel.Frank@zohomail.eu          Haithem Afli Department of Computer Science Munster Technological University Haithem.Afli@mtu.ie
@@ 1 Introduction
Embedding model No augmentation Paraphrasing Summarising
Extracting keywords
Avg. STS22 Avg. STS22 Avg. STS22 Avg. STS22
glove.840B.300d 57.86 54.08 59.98 57.78 60.64 61.21 60.82 56.94
bert-large-cased 60.78 55.48 66.29 58.57 66.89 63.47 62.62 54.64
all-MiniLM-L6-v2 78.91 67.26 80.32 68.47 79.90 69.27 77.96 64.57
all-mpnet-base-v2 80.28 68.00 81.39 69.01 81.48 70.12 79.95 66.19
embed-english-light-v3.0 78.75 67.88 80.30 68.42 80.21 68.97 78.53 67.27
embed-english-v3.0 81.25 68.22 82.10 68.63 82.08 70.29 81.26 66.91
voyage-2 82.50 65.26 83.38 66.80 82.74 68.50 81.21 65.29
voyage-lite-02-instruct 85.95 78.62 85.99 79.08 85.43 79.87 83.91 76.60
voyage-large-2 83.60 63.97 83.92 64.69 83.66 65.90 81.98 63.68
voyage-large-2-instruct 84.61 66.59 84.80 67.72 84.54 68.66 83.40 65.73
mxbai-embed-large-v1 84.63 68.73 85.01 70.03 84.45 70.36 83.42 67.83
llama-3-8b 71.36 33.39 75.28 50.55 76.74 59.44 74.95 51.79
Table 1: Average scores across MTEB STS datasets and STS22 individually for all embedding models with and without generative augmentation using GPT-3.5 Turbo in %, bold: highest average scores for average and STS22 respectively).
Dataset
No Aug-
-mentation
Keyword
Extraction
Random
Keyword
Extraction
STSB 50.73 59.26 23.95
STS12 57.50 58.09 34.78
STS13 70.98 71.12 22.25
STS14 60.69 60.34 28.48
STS15 70.85 72.44 33.86
STS16 63.85 65.31 24.99
STS17 62.05 75.92 36.47
STS22 54.08 56.94 38.44
SICK-R 55.42 55.34 40.04
BIOSSES 32.45 33.50 29.02
Average 57.86 60.82 31.23
Table 2: STS scores per dataset for GloVe embeddings without generative augmentation vs. generative augmentation using LLM-extracted keywords with GPT-3.5 Turbo or random keyword extraction (Spearman’s rank correlation in %, bold: highest scores).
@@
Table 3: Average word count using GPT-3.5 Turbo for paraphrasing, summarising and extracting keywords.
Dataset Paraphrases Summaries
STSB 0.39 0.52
STS12 0.39 0.49
STS13 0.33 0.41
STS14 0.37 0.47
STS15 0.42 0.53
STS16 0.37 0.40
STS17 0.41 0.59
STS22 0.39 0.21
SICK-R 0.46 0.67
BIOSSES 0.46 0.52
Average 0.40 0.48
Table 4: Average Jaccard Similarity between original texts and paraphrases and summaries generated with GPT-3.5 Turbo respectively.
Augmentation
Original text 1 vs
Augmented text 1
Original text 2 vs
Augmented text 2
Augmented text 1 vs
Augmented text 2
Original text 1 vs
Original text 2
Paraphrasing 85.44 85.24 57.31
$\uparrow$
59.81
$\downarrow$
Summarising 85.53 85.36 57.60
Keyword Extraction 84.27 84.20 59.81
Table 5: Average cosine similarity between original text pairs and augmented texts based on all-mpnet-base-v2 embeddings in %. (The last column shows cosine similarity between the original, non-augmented text pairs.)
Representation learning has emerged as a fundamental technique in natural language processing (NLP). However, the quality and robustness of the embeddings are highly dependent on the richness and diversity of the training data. Recent advancements of generative Large Language Models (LLMs) led to remarkable capabilities in generating human-like text. LLMs were also used for data augmentation by Dai et al. (2023), who applied paraphrasing techniques to train a BERT model (Devlin et al., 2019). Wahle et al. (2022) showed that LLM-generated paraphrases are harder to detect for humans than paraphrases generated with simple techniques like synonym replacement. The integration of generative models with representation learning has been explored by augmenting generative models, e.g., in Retrieval-Augmented Generation (Lewis et al., 2020; Guu et al., 2020).
In contrast, we introduce an approach to augment embedding models that applies generative models using test-time compute: Generatively Augmented Sentence Encoding (GASE). Instead of generating synthetic training data, GASE creates textual variants of an input text through paraphrasing, summarising, or extracting keywords at inference time. A joint embedding is derived by pooling the embeddings of the original text and the generated transformation(s). The underlying hypothesis of our work is that adding textual diversity using generative models increases the ability of current embedding models to model semantics and hence benefits the performance of downstream STS tasks.
Also aiming to propose a training-free method, Lei et al. (2024) introduced Meta-Task Prompting, which employs use-case-specific prompts to generate multiple embeddings via a generative model, after which the embeddings are pooled. For retrieval systems, Gao et al. (2023) and Wang et al. (2023) described query expansion approaches using generative models that generate synthetic documents as a response to a retrieval request. Extending this work, GASE combines generative and embedding models without any assumptions on the type of models (unlike Meta-Task Prompting), nor is it limited to augmentation of retrieval queries such as Gao et al. (2023) and Wang et al. (2023). More recently, GenEOL (Thirukovalluru and Dhingra, 2025), a related approach, was introduced (see Table 11 in the Appendix for a comparison with GASE).
2 Method
GASE performs the following steps (see Figure 1):
Generative Augmentation. We apply $k$ different generative models [^1] Using a single generative model for $k=2$ did not yield textually diverse and meaningful variations. to produce $k$ variations of the original input sequence using exactly one of the following transformations: Paraphrasing, which provides a semantically equivalent but lexically or syntactically different variation of the input text. Summarising, which produces a shorter output text that captures the most important information of the input text. Extracting Keywords, which lists the most relevant words from a given text. Figure 2 provides an example for each. We evaluated $k\in\{0,1,2,3\}$ using GPT-3.5 Turbo (OpenAI, 2024b), Reka-Flash (Reka, 2024), and GPT-4o mini  (OpenAI, 2024a) [^2] For model versions and hyperparameters see Appendix F..
2. Sentence Embedding. Generating embeddings for the original and $k$ synthetic texts with one of $12$ encoders. [^3] See Appendix D for a list incl. references.
3. Pooling. The embeddings of the original text and the $k$ synthetic texts are pooled by computing their arithmetic mean. [^4] Mean pooling dominates max pooling (see Table 12).
We evaluate our approach on the English subtasks of the MTEB STS task [^5] See Appendix E for details on the datasets. using cosine similarity and Spearman’s rank correlation [^6] See  Reimers et al. (2016) for why Spearman’s rank correlation is preferable over Pearson correlation for STS tasks..
3 Results
Embedding model No Aug- gpt-3.5-turbo reka-flash gpt-4o-mini gpt-3.5-turbo gpt-3.5-turbo
mentation + reka-flash + reka-flash
+ gpt-4o-mini
glove.840B.300d 57.86 59.98 60.73 60.04 61.97 62.18
bert-large-cased 60.78 66.29 65.76 65.01 68.01 68.41
all-MiniLM-L6-v2 78.91 80.32 79.74 79.81 80.64 80.76
all-mpnet-base-v2 80.28 81.39 81.01 81.07 81.72 81.85
embed-english-light-v3.0 78.75 80.30 79.72 79.90 80.54 80.66
embed-english-v3.0 81.25 82.10 82.00 81.82 82.53 82.08
voyage-2 82.50 83.38 83.37 83.07 83.80 83.72
voyage-lite-02-instruct 85.95 85.99 85.92 85.97 86.05 86.01
voyage-large-2 83.60 83.92 83.98 83.95 84.12 84.17
voyage-large-2-instruct 84.61 84.80 84.68 84.78 84.86 84.91
mxbai-embed-large-v1 84.63 85.01 84.80 84.90 85.17 85.16
llama-3-8b 71.36 75.28 77.02 74.71 77.85 77.52
Table 6: Average STS scores without and with paraphrase augmentation using different generative models in % (bold: highest score per emb. model, underlined: highest score for $k<=1$ per emb. model).
Table 1 shows the results for different augmentation strategies vs. the respective baseline (see Table 13 in the Appendix for the results per dataset). All embedding models improved their avg. scores through augmentation with the greatest improvements for Llama-3 (+6.496.49pp), BERT (+6.116.11pp), and GloVe (+2.962.96pp). Paraphrase augmentation worked best for most models; however, GloVe improved most with keyword extraction, while BERT, MPNET, and Llama-3 excelled with summarisation.
To validate our approach we compared the augmentation with Keyword Extraction using LLMs to randomly extracting keywords [^7] See Appendix B for implementation details. which demonstrated that LLMs extract semantically more meaningful keywords (Table 2). [^8] Additionally, we compared LLM-based keyword extraction with stopword removal showing that it generates semantically richer extractions (see Table 14 in the Appendix).
Summarising and extracting keywords significantly reduced the average word count across datasets (see Table 3). However, GPT-3.5 Turbo could not maintain STS22’s text length when paraphrasing and at the same time only effectively summarised STS22. Comparing the Jaccard Similarity between original texts and paraphrases on the one hand and original texts and summaries on the other hand, Table 4 shows that paraphrases deviate more from the original texts on all datasets except STS22. At the semantic level (quantified by cosine similarity between embeddings) all augmentation strategies preserve meaning: similarities between the transformed and original texts range from $84.20\%$ to $85.36\%$ (Table 5).
@@ 3 Results
Evaluating the different generative models for paraphrase augmentation we find that using a single generative model GPT-3.5 Turbo performed best for all embedding models, except GloVe, Voyage-Large-2, and Llama-3 (Table 6). [^9] For results per dataset see Table 15 and Table 16 in the Appendix).
Using weighted averages between paraphrases and originals peaked at 25%/75% (excl. BERT) and was worst with paraphrases only (Figure 3).
Applying more than one generative model for augmentation yielded higher scores than a single model (see Table 6) with BERT (+7.637.63pp), Llama-3 (+6.496.49pp) and GloVe (+4.324.32pp) showing the largest gains. Moreover, the overall performance of embedding models with lower non-augmented performance was associated with larger performance improvements through generative augmentation.
Embedding model Augment. STS17 STS22
bert-base-multi-
lingual-cased
None 37.16 29.50
Paraphrase 43.93 30.36
Summarise 47.13 42.03
paraphrase-multiling-
ual-mpnet-base-v2
None 82.51 61.56
Paraphrase 82.95 62.72
Summarise 82.88 66.70
multilingual-
e5-large-instruct
None 83.39 69.64
Paraphrase 83.95 66.48
Summarise 83.60 70.26
voyage-multi-
lingual-2
None 85.14 69.88
Paraphrase 85.73 70.21
Summarise 85.41 71.92
Table 7: Average STS scores for multilingual datasets with and without augmentation with GPT-4o mini in % (bold: highest scores per embedding model and dataset).
Generative Model Size (Qwen2.5)
None 0.5b 1.5b 3b 7b 14b
80.28 79.20 79.79 80.26 80.33 80.79
Table 8: Average STS scores of all-mpnet-base-v2 without and with paraphrase augmentation using Qwen2.5 models with different number of parameters in %.
@@
Augmen-
tation
PC IR
glove.840B.300d
None 58.41 11.83
Paraphrase 64.64 15.11 (14.12)
bert-large-cased
None 55.07 4.62
Paraphrase 62.36 6.66 (5.62)
all-MiniLM-L6-v2
None 82.34 31.57
Paraphrase 83.80 33.09 (31.82)
all-mpnet-base-v2
None 83.00 33.29
Paraphrase 84.61 35.21 (34.39)
embed-english-
light-v3.0
None 80.18 21.56
Paraphrase 82.88 24.02 (21.63)
embed-english-
v3.0
None 83.54 35.55
Paraphrase 85.12 35.10 (34.34)
voyage-2
None 83.81 29.65
Paraphrase 85.24 30.34 (28.96)
voyage-lite-02-
instruct
None 93.08 35.95
Paraphrase 92.22 36.38 (35.50)
voyage-large-2
None 92.96 33.20
Paraphrase 92.36 32.83 (31.43)
voyage-large-2-
instruct
None 85.21 37.87
Paraphrase 86.22 37.65 (37.44)
mxbai-embed-
large-v1
None 87.16 37.73
Paraphrase 87.38 38.89 (37.94)
llama-3-8b
None 70.10 25.39
Paraphrase 72.78 28.24 (26.39)
Table 9: Avg. Pair Classification (PC) and Information Retrieval (IR) scores with and without paraphrase augmentation (bold: highest scores per embedding model and task). The bracketed IR score shows augmentation applied only to the queries, not the corpus.
Augmentation all-mpnet-base-v2
None 80.28
paraphrases + summaries
82.26
paraphrases + keywords
82.06
summaries + keywords
82.03
paraphrases + summa-
ries + keywords
82.79
Table 10: Average STS scores for different combinations of augmentations generated with GPT-3.5-Turbo for all-mpnet-base-v2 embeddings in % (bold: highest score).
3.1 Extensions
To evaluate the generalisability of GASE we extended our work as follows:
Multilingual. Consistent with prior findings, GASE evaluated on the $2$ multilingual MTEB STS datasets containing $12$ different languages using $3$ multilingual embedding models yields greater improvements for weaker encoders, while summarisation remains most effective for BERT and STS22 (Table 7). [^10] See Table 17 and Table 18 for results per dataset.
Generative model ablation. Ablating the generative model size of Qwen 2.5  (Yang et al., 2025) shows that GASE’s performance correlates with generative model size (Table 8). [^11] See Appendix F for implementation details.
Additional task. We evaluated GASE on Pair Classification (PC) and Information Retrieval (IR). PC was measured with average precision on Twitter SemEval 2015, Twitter URL Corpus, and Sprint Duplicate Questions, while IR used NDCG@10 on NFCorpus. Consistent with STS, GASE improved scores for most models, with gains negatively correlated with baseline performance (Table 9). [^12] See Table 19 in the Appendix for scores per dataset and Muennighoff et al. (2023) for details on score calculations.
Cross-augmentation. Rather than applying a single augmentation strategy across $k$ different generative models, we used a single generative model with multiple augmentation strategies. Tested with all-mpnet-base-v2, using $3$ augmentations the approach outperformed all other (Table 10).